OpenStudy (anonymous):

write an equation of the line that passes through the given points (3/2,-1/2) and is perpendicular to y=-1/4x+11

OpenStudy (anonymous):

the slope of \(y=-\frac{1}{4}x+11\) is ?

OpenStudy (anonymous):

is it clear or not? if not, i will tell you

OpenStudy (anonymous):

i know i have t o plug in the two fractions (3/2 and -1/2) into the y=mx+b form with slope slope of positive 4.

OpenStudy (anonymous):

yeah! good